OpenStudy (kaydra14):

Kailey plays video games for x hours a month. Trent plays video games four hours more than Kailey and Megan plays two times more than Trent. Which expression represents the amount of time Megan spends playing video games? A)2(4x) B)2x + 4 C)2x - 4 D)2(x + 4)

OpenStudy (anonymous):

What do you guess?

OpenStudy (kaydra14):

a

OpenStudy (anonymous):

-.-

OpenStudy (texaschic101):

Kailey plays for x hrs Trent plays 4 hrs more then Kailey....x + 4 Megan plays 2 times more then trent .... 2(x + 4)

OpenStudy (texaschic101):

do you see how I arrived at that ? Do you have any questions at all ?

OpenStudy (kaydra14):

yes, i understand thank you so much

OpenStudy (texaschic101):

anytime...glad to help. With problems such as these, you have to basically break them down to solve them :)

OpenStudy (kaydra14):

can you please help me with a next question please

OpenStudy (texaschic101):

I can try....but I can't make any promises that I will know how to do it.

OpenStudy (texaschic101):

ask away :)

OpenStudy (kaydra14):

Mrs. Reed is decorating wreaths. She needs 24 in. of gold ribbon and 45 in. of white ribbon to tie around each wreath. The white ribbon is sold by the foot. About how many feet of white ribbon will Mrs. Reed need to make 18 wreaths? A)30 ft B)50 ft C)70 ft. D)90 ft.

OpenStudy (texaschic101):

45 inches of white ribbon for each wreath.....and she needs 18 wreaths so....45 x 18 = 810....she needs 810 inches of white ribbon. However, we need this in feet. There are 12 inches in 1 ft. So we divide 810 by 12.....810/12 = 67.5 ft. Closest answer is 70 ft. questions ?

OpenStudy (kaydra14):

thank you , now i saw where i went wrong

OpenStudy (texaschic101):

sometimes word problems throw in extra info into the problem that is not even needed. That can mess with someone's brain...lol

OpenStudy (kaydra14):

Charlie's weight is nine kilograms greater than his brother's. Together they weigh 99 kilograms. Let x represent the weight of the little brother. Which equation can be used to determine what each one weighs? A)x2 + 9 = 99 B)2x - 9 = 99 C)2x + 9 = 99 D3x - 9 = 99